OpenStudy (anonymous):

((sinx+cosx)^2−1)/sinxcosx=2 Only work on one side

OpenStudy (anonymous):

= (sin^2 x +2sinx cosx + cos^2x +1) / sinxcosx = 1-1 +2sinxcosx /sinxcosx =2

OpenStudy (anonymous):

is that okay with you?

OpenStudy (anonymous):

yes that's great, thank you